Essay: Must SHOW ALL WORK! You want to create a rectangular garden, its length is 3x-3y ft. and x-5y ft. in width. Be sure to multiply this problem out class, and express it in simplest Correct mathematical form including units.
_______________________________________________________________________________
Thank you in advance.

Found 2 solutions by solver91311, dfrazzetto:
Answer by solver91311(24713)



You haven't asked a question. What do you need to know? What is it that you want to calculate? The area? The perimeter?

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by dfrazzetto(283)

What do you want? The Area? Perimeter
Area will be in ft^2, perimeter in ft
Area:
A = LxW = (3x-3y)(x-5y) = 3x^2 -15xy -3xy + 15y^2
A = (3x^2 -18xy + 15y^2) ft^2
Perimeter:
P = 2L + 2W = 2(3x-3y) + 2(x-5y) = 6x - 6y + 2x - 10y
P = (8x - 16y) ft